Shift Summary for Owl Shift, 12/07/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Tue, 7 Dec 1999 06:58:17 -0600

Shift Summary for Owl shift, 12/07/99

On Shift: Ashkan Alavi

========================================================================

Runs ended this shift:

15146 -- COSMICS to disk --

========================================================================

Overview of shift:

Continuation of TeV downtime for magnet repairs.

DAQ running smoothly with COSMICS configuration until ktev2 died around
6:00 am.

'laser_rms too wide' message all night.

Walk-throughs and gas check done.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shifts

Can not connect to ktev2. I think ktev2 should be rebooted.
Keep an eye on CsI channel 2446